# Long-form music generation with latent diffusion

### v1

@@ Abstract
Audio-based generative models for music have seen great strides recently, but so far have not managed to produce full-length music tracks with coherent musical structure. We show that by training a generative model on long temporal contexts it is possible to produce long-form music of up to 4m 45s. Our model consists of a diffusion-transformer operating on a highly downsampled continuous latent representation (latent rate of 21.5 Hz). It obtains state-of-the-art generations according to metrics on audio quality and prompt alignment, and subjective tests reveal that it produces full-length music with coherent structure.
@@ 2.1 Autoencoder
We train this structure using a variety of objectives. Firstly there is a reconstruction loss, consisting of a perceptually weighted multi-resolution STFT loss [28] that deals with stereo signals as follows: the STFT loss is applied to the mid-side (M/S) representation of the stereo audio, as well as the left and right channels separately. This second component is down-weighted by 0.5 compared to the M/S component, and exists to mitigate potential ambiguity around left-right placement. Secondly we employ an adversarial loss term, utilizing a convolutional discriminator model [22]. The discriminator has hyperparameters consistent with previous similar work [14], but with channel count scaled to give approximately 4x the parameter count.
@@ 3 Training Setup
The DiT is trained to predict a noise increment from noised ground-truth latents, following the v-objective [37] approach. We sample from the trained model using DPM-Solver++ [38], with classifier-free guidance [39].
@@ 4.1 Quantitative evaluation
Table 2: Song Describer Dataset (no-singing): results of the 3m 10s pre-trained model. Our pre-trained model generates 3m 10s outputs, but during inference it can generate 2m outputs by relying on the timing conditioning. We trim audios to 2m (discarding the end silent part) for a fair quantitative evaluation against the state-of-the-art (see Section 2.3).
@@
Table 3: Song Describer Dataset (no-singing): results of the 4m 45s fully-trained model. Our fully-trained model generates 4m 45s outputs, but during inference it can generate 2m outputs by relying on the timing conditioning. We trim audios to 2m (discarding the end silent part) for a fair quantitative evaluation against the state-of-the-art (see Section 2.3).
@@ 4.2 Qualitative evaluation
Table 5: Autoencoder reconstructions on the Song Describer Dataset (no-singing). Although different autoencoders operate at various sampling rates, the evaluations are run at 44.1kHz bandwidth for a fair comparison. Sorted by latent rate.
@@ 4.3 Autoencoder evaluation
We evaluate the audio reconstruction quality of our autoencoder in isolation, as it provides a quality ceiling for our system. We achieve this by comparing ground-truth and reconstructed audio via a number of established audio quality metrics[23, 22]: STFT distance, MEL distance and SI-SDR (as in AuraLoss library [28], with its default parameters). The reconstructed audio is obtained by encoding-decoding the ground-truth audio from the Song Describer Dataset (no-singing) through the autoencoder. As a comparison, we calculate the same metrics on a number of publicly available neural audio codecs including Encodec [22], DAC [23] and AudioGen [24]. Encodec and DAC are selected because are widely used for generative music modeling [8, 4, 10]. We select the Encodec 32kHz variant because our MusicGen baseline relies on it, and DAC 44.1kHz because its alternatives operate at 24kHz and 16kHz. Further, since our autoencoder relies on a continuous latent, we also compare against AudioGen, a state-of-the-art autoencoder with a continuous latent. Notably, AudioGen presents both continuous and discrete options, and we report both for completeness. All neural audio codecs are stereo, except DAC 44.1kHz and Encodec 32 kHz. In those cases, we independently project left and right channels and reconstruct from those.

### v2

@@ Abstract
Audio-based generative models for music have seen great strides recently, but so far have not managed to produce full-length music tracks with coherent musical structure from text prompts. We show that by training a generative model on long temporal contexts it is possible to produce long-form music of up to 4m 45s. Our model consists of a diffusion-transformer operating on a highly downsampled continuous latent representation (latent rate of 21.5 Hz). It obtains state-of-the-art generations according to metrics on audio quality and prompt alignment, and subjective tests reveal that it produces full-length music with coherent structure.
@@ 2.1 Autoencoder
We train it using a variety of objectives. First, the reconstruction loss, consisting of a perceptually weighted multi-resolution STFT [28] that deals with stereo audio as follows: the STFT loss is applied to the mid-side (M/S) representation of the stereo audio, as well as the left and right (L/R) channels separately. The L/R component is weighted by 0.5 compared to the M/S one, and exists to mitigate potential ambiguity around L/R placement. Second, an adversarial loss term with feature matching, utilizing 5 convolutional discriminators [22] with hyperparameters consistent with previous work [14], but with channel count scaled to give $\approx$4 times the parameter count. And third, the KL divergence loss term that is weighted by ×$\times 10^{-4}$.
@@ 3 Training Setup
The DiT is trained to predict a noise increment from noised ground-truth latents, following the v-objective [37]. We sample from our model using DPM-Solver++ [38] (100 steps), with classifier-free guidance [39] (scale of 7.0).
@@ 4.1 Quantitative evaluation
Table 2: Song Describer Dataset (no-singing subset): results of the 3m 10s pre-trained model. Our pre-trained model generates 3m 10s outputs, but during inference it can generate 2m outputs by relying on the timing conditioning. We trim audios to 2m (discarding the end silent part) for a fair quantitative evaluation against the state-of-the-art (see Section 2.3).
@@
Table 3: Song Describer Dataset (no-singing subset): results of the 4m 45s fully-trained model. Our fully-trained model generates 4m 45s outputs, but during inference it can generate 2m outputs by relying on the timing conditioning. We trim audios to 2m (discarding the end silent part) for a fair quantitative evaluation against the state-of-the-art (see Section 2.3).
@@ 4.2 Qualitative evaluation
Table 5: Autoencoder reconstructions on the Song Describer Dataset (all the dataset). Although different autoencoders operate at various sampling rates, the evaluations are run at 44.1kHz bandwidth for a fair comparison. Sorted by latent rate.
@@ 4.3 Autoencoder evaluation
We evaluate the audio reconstruction quality of our autoencoder in isolation, as it provides a quality ceiling for our system. We achieve this by comparing ground-truth and reconstructed audio via a number of established audio quality metrics[23, 22]: STFT distance, MEL distance and SI-SDR (as in AuraLoss library [28], with its default parameters). The reconstructed audio is obtained by encoding-decoding the ground-truth audio from the Song Describer Dataset (all the dataset, 706 tracks) through the autoencoder. As a comparison, we calculate the same metrics on a number of publicly available neural audio codecs including Encodec [22], DAC [23] and AudioGen [24]. Encodec and DAC are selected because are widely used for generative music modeling [8, 4, 10]. We select the Encodec 32kHz variant because our MusicGen baseline relies on it, and DAC 44.1kHz because its alternatives operate at 24kHz and 16kHz. Further, since our autoencoder relies on a continuous latent, we also compare against AudioGen, a state-of-the-art autoencoder with a continuous latent. Notably, AudioGen presents both continuous and discrete options, and we report both for completeness. All neural audio codecs are stereo, except DAC 44.1kHz and Encodec 32 kHz. In those cases, we independently project left and right channels and reconstruct from those.